Udon Inc. adopted dollar-value LIFO (DVL) as of January 1, 2021, when it had an inventory of $705,000. Its inventory as of December 31, 2021, was $779,100 at year-end costs and the cost index was 1.06. What was DVL inventory on December 31, 2021?

A. $747,300.

B. $779,100.

C. $735,000.

D. $736,800.
Answer #1

First of all, we need to compute the value of ending inventory at base-year-prices. It is computed using the following formula:

Ending inventory at base-year-prices = 779100/1.06

= 735000

Now we can compute the real-dollar quantity increase in inventory:

(735000- 705000) = 20000

The next step is to value this real dollar quantity increase in inventory at year-end-prices

=(735000- 705000)*1.06 = 21200

DVL inventory on Dec 31 = 705000+21200 = 726200
Answer #2

To calculate the DVL (dollar-value LIFO) inventory on December 31, 2021, we need to determine the inventory value at base-year costs and then adjust it to year-end costs using the cost index.

Given information: Inventory on January 1, 2021 (base-year costs) = $705,000 Inventory on December 31, 2021 (year-end costs) = $779,100 Cost index = 1.06

To find the DVL inventory on December 31, 2021, we'll follow these steps:

  1. Convert the year-end inventory value to base-year costs: DVL Base-year inventory on December 31, 2021 = Year-end inventory / Cost index DVL Base-year inventory on December 31, 2021 = $779,100 / 1.06 DVL Base-year inventory on December 31, 2021 ≈ $734,906.60

  2. Determine the increase in inventory value during the year: Increase in DVL Base-year inventory = DVL Base-year inventory on December 31, 2021 - Inventory on January 1, 2021 Increase in DVL Base-year inventory = $734,906.60 - $705,000 Increase in DVL Base-year inventory ≈ $29,906.60

  3. Calculate the DVL inventory on December 31, 2021: DVL inventory on December 31, 2021 = Inventory on January 1, 2021 + Increase in DVL Base-year inventory DVL inventory on December 31, 2021 ≈ $705,000 + $29,906.60 DVL inventory on December 31, 2021 ≈ $734,906.60

Therefore, the DVL inventory on December 31, 2021, is approximately $734,906.60. However, none of the provided answer choices match this value exactly.


answered by: mervetokaz